OpenStudy (anonymous):

Find dy/dx for y=3x^2+sqrtx/x

OpenStudy (anonymous):

Separate it into two pieces, remember, sqrt(x) = x^(1/2) and x^(1/2)/x = x^(-1/2). You now have two simple pieces that you can derivatize.

OpenStudy (anonymous):

 Am I right so far?